# Uncertainty Quantification of Surrogate Models using Conformal Prediction

### v2

@@ 3.7.1 Model
We apply CP to the Graph-FM model of Oskarsson et al. (2024). Graph-FM is a graph-based neural weather prediction model (Keisler, 2022; Lam et al., 2023), where a hierarchical Graph Neural Network (GNN) is utilized for producing the forecast. Let $X^{t}$ denote the full weather state at time step $t$, including multiple atmospheric variables modeled for all grid cells in some discretization. Examples of such atmospheric variables are temperature, wind, geopotential and solar radiation. The GNN $g$ in Graph-FM represents the single time step prediction
@@
where $F^{t+1}$ are known forcing inputs that should not be predicted. Taking the two past states as inputs to $g$ allows the model to make use of both magnitude and first derivative information. Equation 5 can be applied iteratively to roll out a complete forecast of $T$ time steps. The full forecasting model can thus be viewed as mapping from initial weather states $X^{-1:0}$ and forcing $F^{1:T}$ to a forecast $X^{1:T}$. The forecast $X^{1:T}$ is a tensor of shape $T\times N_{x}\times N_{y}\times N_{\text{var.}}$ where $N_{\text{var.}}$ is the number of atmospheric variables modeled. We consider two versions of Graph-FM, trained with different loss functions:
@@
  • Graph-FM (NLL): A version of Graph-FM that outputs both the mean and standard deviation for each time, variable and grid cell. This model was trained with a Negative Log-Likelihood (NLL) loss, assuming a diagonal Gaussian predictive distribution (also referred to as the uncertainty loss (Chen et al., 2023a)). Apart from the change of loss function the training setup was identical.
For the Graph-FM (MSE) we compute non-conformity scores using the RES strategy. As the Graph-FM (NLL) is probabilistic, we there use STD non-conformity scores. Note that these are computed based on the standard-deviations directly output from the model, rather than from sample-estimates based on MC dropout.
@@ 3.7.2 Limited Area Forecasting
In this first experiment we apply CP to a limited area version of Graph-FM. Forecasts are here produced for a limited area covering the Nordic region. These Graph-FM models were trained on the limited area dataset from Oskarsson et al. (2024), consisting of forecasts from the MEPS system (Müller et al., 2017). One such forecast includes $N_{\text{var.}}=17$ variables modeled on a $N_{x}\times N_{y}=238\times 268$ grid over $T=19\times 3h$ time steps (up to 57 hour lead time). When Graph-FM is used in a limited area configuration, it produces weather forecasts for a specific sub-area of the globe. To achieve this, boundary conditions along the edges of the forecasting area are given as important forcing inputs. The exact models used have 4 graph processing layers and use 64-dimensional latent representations. We refer to Oskarsson et al. (2024) for further details about the model and data.
We use forecasts started during September 2021333For calibration we specifically use forecasts started during the dates 2021-09-04 – 2021-09-30. The model was trained on forecasts started during the last days of August, which are rolled out over the first days of September. To avoid strong correlations to the training data we use only forecasts from September 4 onwards for calibration. as our calibration data set and forecasts started during September 2022 as test data. By using the same month for calibration and testing we minimize the effect of distributional shifts due to seasonal effects. Having access to calibration data from the same month, collected the previous year, is a reasonable assumption in practical settings.
@@
Figure 14 shows the empirical coverage for the test set. With CP we can achieve calibrated uncertainty estimates for both versions of Graph-FM. Of great interest in the weather forecasting setting is the uncertainty for specific future time points. We visualise this by plotting the width of the error bars for all spatial locations at different lead times in an example forecast. Such plots for shortwave solar radiation are shown in Figure 15 and for geopotential in Figure 16.
Figure 15 highlights an important difference between the two methods for computing non-conformity scores. As the shortwave solar radiation is close to 0 during the night, it is easy for the model to predict. During the day this is far more challenging. With the RES non-conformity scores, used for Graph-FM (MSE) in Figure 15(a), the width of the predictive intervals are determined during calibration, and does not change depending on the forecast from the model. As a specific lead time can fall both during day and night, depending on the initialization time, CP will give large error bars also during the night. This can be compared to the results for Graph-FM (NLL) in Figure 15(b), using STD non-conformity scores. In this case the bounds are very tight for lead times during the night (33 h and 57 h). It can also be noted that for Graph-FM (NLL) at lead time 15 we see clear spatial features appearing in the error bars themselves. This corresponds to higher forecast uncertainty in areas of rapid change. The forecast-dependent patterns for Graph-FM (NLL) thus have desirable properties, but this relies on having a model that outputs (potentially uncalibrated) standard-deviations.
@@
Figure 15: Width of predictive interval at $\alpha=0.05$ for shortwave solar radiation (nswrs) in an example forecasts. Note that for Graph-FM (MSE) these widths are constant after calibration, while for Graph-FM (NLL) they depend on the predicted standard deviations for the specific forecast. This is most noticeable by Graph-FM (NLL) having very tight bounds during the night (lead times 33 h and 57 h), when the short wave solar radiation is close to 0 and easy to predict.
@@ 3.7.3 Global Forecasting
We next experiment with CP for global weather forecasting. The models used are again Graph-FM (MSE) and Graph-FM (NLL), but here applied on the full globe. These models are trained on a version of the ERA5 reanalysis dataset (Hersbach et al., 2020) using a 1.5° latitude-longitude grid. The global models have 8 graph processing layers and use 256-dimensional latent representations. Each forecast includes 5 surface variables and 6 atmospheric variables, each modeled at 13 different vertical pressure levels in the atmosphere. Due to the large number of variables forecast (83 in total) we here only perform CP for a subset of these. This subset includes all surface variables and the atmospheric variables at pressure level 700 hPa. This results in a total of $N_{\text{var.}}=11$ variables, modeled on a $N_{x}\times N_{y}=240\times 121$ grid over $T=40$ time steps (up to 10 days lead time with 6 h time steps). We note that a strength of the CP framework is that uncertainty quantification can be performed per variable, alleviating memory issues during calibration. Because of this the procedure could trivially be extended to the full set of variables, as long as the full forecasts are stored. We again refer to Oskarsson et al. (2024) for more details about the global models and data configuration.
For the global experiment we use full years of forecasts for calibration and evaluation, all started from ERA5 as initial conditions. Forecasts at 00 and 12 UTC each day of 2018 are used as the calibration set and a similar set of forecasts for 2019 used as test set [^4] We remove forecasts started during the last 10 days of 2018 from the calibration set to avoid strong correlations with the test data at the start of 2019. Note that forecasts in the test set started during the last days of 2019 will extend into time points in 2020.. The ground truth is given by ERA5 at each forecasted time point. Using a full year for calibration allows for capturing the model performance across all different seasons. This allows for calibrating the model once, and then using the computed $\hat{q}$ values for the full next year of forecasts. However, any distributional shift due to climate variations from one year to the next remain. We note that in practice this does not seem to cause any major issue for achieving the desired coverage.
Figure 17 shows an example prediction from Graph-FM (MSE) and corresponding error bars. Global forecasting up to 10 days is a more challenging task than the limited area modeling up to 57 h. We see that at 10 days the model prediction fails to capture much of the patterns in the ground truth data. Importantly, this is accurately captured in the error bars, which increase with the lead time to high values at 10 days.
@@
As for all experiments, we include results for the weather forecasting models in Table 1. In both the global and limited area setting CP successfully produces calibrated error bars. For the Graph-FM (NLL) model the original standard deviations output by the model are too low, leading to invalid error bars and insufficient coverage. After applying CP however the error bars are nicely calibrated. We generally see that the Graph-FM (NLL) has tighter error bars than Graph-FM (MSE). This can be attributed to these being input-dependent, specific to each forecast from the model.
@@ 3.7.4 Discussion on Exchangeability
Figure 20: Constructing exchangeable input-output pairs from a time series dataset. Calibration dataset is arbitrarily characterised by a spatio-temporal dataset spanning 100 time instances, where the model takes in the field at a day and autoregressively predicts for the next 3 time instances. Since the model represents an initial boundary value problem, we are able to break each input-output ($X_{i},Y_{i}$) as an exchangeable data point, essentially sampled from a distribution spanning the entire month.
@@
Traditionally the CP framework is limited in application to time-series modelling as it fails the exchangeability assumption. Previous research has looked into fixing this violation of exchangeability by accounting for the distribution shift using weighted conformal techniques (Tibshirani et al., 2019), but becomes limited in application in multi-variate settings. Other work has explored CP for multi-variate time series forecasting, where each time-series is treated as an exchangeable observation (Stankeviciute et al., 2021). Within the weather modelling tasks outlined in this section, we maintain exchangeability by treating each modelling task as an Initial Boundary Value Problem (IBVP, boundary given by the forcing terms in Equation 5). As given in Equation 5, the model takes in the initial conditions, $\mathcal{X}\in\mathbb{R}^{T_{\text{in}=1}\times N_{x}\times N_{y}\times N_{var}}$ and is auto-regressively rolled out $T_{out}$ steps to obtain the output $\mathcal{Y}\in\mathbb{R}^{T_{\text{out}}\times N_{x}\times N_{y}\times N_{var}}$. Being dependent on the initial conditions alone and being rolled out for a fixed number of steps, each input-output pair as mathematically outlined in Section 2.1.2 and visually represented in Figure 20 can be treated as an exchangeable pair. We are allowed to make this assumption on exchangeability since the model is agnostic to the temporal nature of the dataset beyond the autoregressive roll-out of each forward prediction, typical of an initial boundary value problem i.e. the neural weather forecast starting from 18:00 $3^{rd}$ January is independent of the forecast made using the neural weather models starting at 12:00 $1^{st}$ January. Here the calibration dataset is seen as samples from an extremely large distribution which effectively characterises the entirety of the entire month/year under consideration. Thus, by combining our preservation of spatio-temporal structure as outlined in Section 2.1.2 and by treating the neural weather models as initial boundary value problems, we are able to maintain exchangeability across the calibration datasets, allowing us to perform CP.
@@
Though the above description discusses about exchangeability across the calibration dataset, it does not extend across to the prediction set. For each of the experiments within the limited area and global weather forecasting, we make the assumption that the climate does not vary significantly across the years under consideration for the calibration and prediction sets.
@@ 3.8 Camera Diagnostic on a Tokamak
The Mega-Ampere Spherical Tokamak (MAST) is an experimental fusion device at the UK Atomic Energy Authority. It was equipped with fast camera imaging diagnostics at several points across the Tokamak for capturing images within the visible spectrum, capturing the plasma evolution in real-time  [^5] https://ccfe.ukaea.uk/research/mast-upgrade/. These Photron cameras  [^6] https://photron.com/ are essential diagnostics that contributed to fundamental understandings of key plasma phenomena in Tokamaks in recent years (Kirk et al., 2006). While camera data have mostly been used for qualitative analysis, recently it have been exploited to provide statistical insight into plasma turbulence (Walkden et al., 2022), as well as global instabilities (disruptions) (Ham et al., 2022).
In our previous work (Gopakumar et al., 2024), we demonstrated the utility of the FNO to forecast the evolution of the plasma within the Tokamak. Within that work, the FNO was trained to take in an input sequence of 10 camera frames and predict a sequence of the next 10 camera frames of the plasma within the Tokamak. The model is trained on sequences of frames characterising the entire duration of a plasma shot within the MAST Tokamak. In this work, we apply the CP framework across the FNO-camera using the AER nonconformity score to provide valid error bars to the plasma evolution forecasts. For results and further information about the camera data, FNO and the training setup used, we advise the reader to refer to the original work by (Gopakumar et al., 2024).
Though the FNO is trained to predict the evolution of plasma for the entire time duration of a plasma shot, similar to Section 3.7, the modelling task and the data are structured to represent that of an initial value problem. As represented in Figure 20, the model takes in a fixed time sequence and outputs a fixed time sequence, its output is fully determined by the initial sequence of frames provided to the model. In our framework, we apply CP across the entire spatio-temporal domain of the output simultaneously and not sequentially, where the temporal nature of the output (the 10 time frames) is kept intact. Figure 22 visualises the error bars obtained using the CP framework over the FNO used to forecast the plasma evolution across the central solenoid of the Tokamak and obtains exact coverage as indicated in Appendix I.
We have the following arguments to support the validity of using CP for this time dependant problem, 1) only the calibration dataset and the prediction $X_{n+1}$ are required to be exchangeable, 2) we predict errors over an entire time interval simultaneously, not sequentially, and 3) we assume there is no large shift in the data between the calibration shot and prediction shot. The exchangeability requirement across the calibration and prediction shot requires them to follow similar plasma discharge profiles and device conditions. Predicting across a characteristically different shot violates this assumption and suffers from lack of coverage as shown in Figure 21.
@@
Figure 21: Comparing coverage obtained across prediction shots that are exchangeable (similar plasma profiles) to the calibration shot and that is in-exchangeable (dissimilar plasma profiles.
@@
Figure 22: Camera (top), FNO (middle) and the prediction interval width obtained using CP with $\alpha=0.5$ (bottom).

### v3

@@ 3.7.1 Model
We apply CP to the Graph-FM model of neural_lam. Graph-FM is a graph-based neural weather prediction model (keisler; lam2022graphcast), where a hierarchical Graph Neural Network (GNN) is utilised for producing the forecast. Let $X^{t}$ denote the full weather state at time step $t$, including multiple atmospheric variables modelled for all grid cells in some discretisation. Examples of such atmospheric variables are temperature, wind, geopotential and solar radiation. The GNN $g$ in Graph-FM represents the single time step prediction
@@
where $F^{t+1}$ are known forcing inputs that should not be predicted. Taking the two past states as inputs to $g$ allows the model to make use of both magnitude and first derivative information. Equation 5 can be applied iteratively to roll out a complete forecast of $T$ time steps. The full forecasting model can thus be viewed as a mapping from initial weather states $X^{-1:0}$ and forcing $F^{1:T}$ to a forecast $X^{1:T}$. The forecast $X^{1:T}$ is a tensor of shape $T\times N_{x}\times N_{y}\times N_{\text{var.}}$ where $N_{\text{var.}}$ is the number of atmospheric variables modelled. We consider two versions of Graph-FM, trained with different loss functions:
@@
  • Graph-FM (NLL): A version of Graph-FM that outputs both the mean and standard deviation for each time, variable and grid cell. This model was trained with a Negative Log-Likelihood (NLL) loss, assuming a diagonal Gaussian predictive distribution (also referred to as the uncertainty loss (fengwu)). Apart from the change of loss function, the training setup was identical.
For the Graph-FM (MSE) we compute non-conformity scores using the AER strategy. As the Graph-FM (NLL) is probabilistic, we use STD non-conformity scores. Note that these are computed based on the standard deviations directly output from the model, rather than from sample estimates based on MC dropout.
@@ 3.7.2 Limited Area Forecasting
In this first experiment, we apply CP to a limited area version of Graph-FM. Forecasts are here produced for a limited area covering the Nordic region. These Graph-FM models were trained on the limited area dataset from neural_lam, consisting of forecasts from the MEPS system (meps). One such forecast includes $N_{\text{var.}}=17$ variables modelled on a $N_{x}\times N_{y}=238\times 268$ grid over $T=19\times 3h$ time steps (up to 57-hour lead time). When Graph-FM is used in a limited area configuration, it produces weather forecasts for a specific sub-area of the globe. To achieve this, boundary conditions along the edges of the forecasting area are given as important forcing inputs. The exact models used have 4 graph processing layers and use 64-dimensional latent representations. We refer to neural_lam for further details about the model and data.
We use forecasts started during September 2021111For calibration, we specifically use forecasts started during the dates 2021-09-04 – 2021-09-30. The model was trained on forecasts started during the last days of August, which are rolled out over the first days of September. To avoid strong correlations to the training data, we use only forecasts from September 4 onwards for calibration. As our calibration data set forecasts that started during September 2022 as test data. By using the same month for calibration and testing, we minimise the effect of distributional shifts due to seasonal effects. Having access to calibration data from the same month, collected the previous year, is a reasonable assumption in practical settings.
@@
Figure 14 shows the empirical coverage for the test set. With CP, we can achieve calibrated uncertainty estimates for both versions of Graph-FM. Of great interest in the weather forecasting setting is the uncertainty for specific future time points. We visualise this by plotting the width of the error bars for all spatial locations at different lead times in an example forecast. Such plots for shortwave solar radiation are shown in Figure 15 and for geopotential in Figure 16.
Figure 15 highlights an important difference between the two methods for computing non-conformity scores. As the shortwave solar radiation is close to 0 during the night, it is easy for the model to predict. During the day, this is far more challenging. With the AER non-conformity scores, used for Graph-FM (MSE) in Figure 15(a), the width of the predictive intervals is determined during calibration, and does not change depending on the forecast from the model. As a specific lead time can fall both during the day and night, depending on the initialisation time, CP will give large error bars also during the night. This can be compared to the results for Graph-FM (NLL) in Figure 15(b), using STD non-conformity scores. In this case, the bounds are very tight for lead times during the night (33 h and 57 h). It can also be noted that for Graph-FM (NLL) at lead time 15, we see clear spatial features appearing in the error bars themselves. This corresponds to higher forecast uncertainty in areas of rapid change. The conditional dependency that emerges while using STD in Graph-FM (NLL) thus has desirable properties, but this relies on having a model that outputs (potentially uncalibrated) standard deviations.
@@
Figure 15: Width of predictive interval at $\alpha=0.05$ for shortwave solar radiation (nswrs) in an example forecasts. Note that for Graph-FM (MSE) these widths are constant after calibration, while for Graph-FM (NLL) they depend on the predicted standard deviations for the specific forecast. This is most noticeable by Graph-FM (NLL) having very tight bounds during the night (lead times 33 h and 57 h), when the short-wave solar radiation is close to 0 and easy to predict.
@@ 3.7.3 Global Forecasting
We next experiment with CP for global weather forecasting. The models used are again Graph-FM (MSE) and Graph-FM (NLL), but here applied on the full globe. These models are trained on a version of the ERA5 reanalysis dataset (era5) using a 1.5° latitude-longitude grid. The global models have 8 graph processing layers and use 256-dimensional latent representations. Each forecast includes 5 surface variables and 6 atmospheric variables, each modelled at 13 different vertical pressure levels in the atmosphere. Due to the large number of variables forecast (83 in total), we here only perform CP for a subset of these. This subset includes all surface variables and the atmospheric variables at pressure level 700 hPa. This results in a total of $N_{\text{var.}}=11$ variables, modelled on a $N_{x}\times N_{y}=240\times 121$ grid over $T=40$ time steps (up to 10 days lead time with 6 h time steps). We note that a strength of the CP framework is that uncertainty quantification can be performed per variable, alleviating memory issues during calibration. Therefore, the procedure could trivially be extended to the full set of variables, as long as the full forecasts are stored. We again refer to neural_lam for more details about the global models and data configuration.
For the global experiment, we use full years of forecasts for calibration and evaluation, all starting from ERA5 as initial conditions. Forecasts at 00 and 12 UTC each day of 2018 are used as the calibration set, and a similar set of forecasts for 2019 is used as the test set [^2] We remove forecasts started during the last 10 days of 2018 from the calibration set to avoid strong correlations with the test data at the start of 2019. Note that forecasts in the test set started during the last days of 2019 will extend into time points in 2020.. The ground truth is given by ERA5 at each forecasted time point. Using a full year for calibration allows for capturing the model performance across all different seasons. This allows for calibrating the model once, and then using the computed $\hat{q}$ values for the full next year of forecasts. However, any distributional shift due to climate variations from one year to the next remains. We note that in practic,e this does not seem to cause any major issue for achieving the desired coverage.
Figure 17 shows an example prediction from Graph-FM (MSE) and corresponding error bars. Global forecasting up to 10 days is a more challenging task than the limited area modelling up to 57 h. We see that at 10 days the model prediction fails to capture much of the patterns in the ground truth data. Importantly, this is accurately captured in the error bars, which increase with the lead time to high values at 10 days.
@@
As for all experiments, we include results for the weather forecasting models in Table 1. In both the global and limited area setting CP successfully produces calibrated error bars. For the Graph-FM (NLL) model the original standard deviations output by the model are too low, leading to invalid error bars and insufficient coverage. After applying CP however the error bars are well calibrated. We generally see that the Graph-FM (NLL) has tighter error bars than Graph-FM (MSE). This can be attributed to these being input-dependent, specific to each forecast from the model.
@@ 3.7.4 Discussion on Exchangeability
Figure 20: Constructing exchangeable input-output pairs from a time series dataset. The calibration dataset is arbitrarily characterised by a spatio-temporal dataset spanning 100 time instances, where the model takes in the field at a day and autoregressively predicts for the next 3 time instances. Since the model represents an initial boundary value problem, we are able to break each input-output ($X_{i},Y_{i}$) as an exchangeable data point, essentially sampled from a distribution spanning the entire month.
@@
Traditionally, the CP framework is limited in application to time-series modelling as it fails the exchangeability assumption. Previous research has looked into fixing this violation of exchangeability by accounting for the distribution shift using weighted conformal techniques (tibshirani2019conformal), but becomes limited in application in multi-variate settings. Other work has explored CP for multi-variate time series forecasting, where each time-series is treated as an exchangeable observation (conformaltimeserires). Within the weather modelling tasks outlined in this section, we maintain exchangeability by treating each modelling task as an Initial Boundary Value Problem (IBVP, boundary given by the forcing terms in Equation 5). As given in Equation 5, the model takes in the initial conditions, $\mathcal{X}\in\mathbb{R}^{T_{\text{in}=1}\times N_{x}\times N_{y}\times N_{var}}$ and is auto-regressively rolled out $T_{out}$ steps to obtain the output $\mathcal{Y}\in\mathbb{R}^{T_{\text{out}}\times N_{x}\times N_{y}\times N_{var}}$. Being dependent on the initial conditions alone and being rolled out for a fixed number of steps, each input-output pair as mathematically outlined in Section 2.1.1 and visually represented in Figure 20 can be treated as an exchangeable pair. We are allowed to make this assumption on exchangeability since the model is agnostic to the temporal nature of the dataset beyond the autoregressive roll-out of each forward prediction, typical of an initial boundary value problem i.e. the neural weather forecast starting from 18:00 $3^{rd}$ January is independent of the forecast made using the neural weather models starting at 12:00 $1^{st}$ January. Here the calibration dataset is seen as samples from an extremely large distribution which effectively characterises the entirety of the entire month/year under consideration. Thus, by combining our preservation of spatio-temporal structure as outlined in Section 2.1.1 and by treating the neural weather models as initial boundary value problems, we are able to maintain exchangeability across the calibration datasets, allowing us to perform CP.
@@
Though the above description discusses about exchangeability across the calibration dataset, it does not extend across to the prediction set. For each of the experiments within the limited area and global weather forecasting, we assume that the climate does not vary significantly across the years under consideration for the calibration and prediction sets.
@@ 3.8 Camera Diagnostic on a Tokamak
The Mega-Ampere Spherical Tokamak (MAST) at the UK Atomic Energy Authority was equipped with fast Photron camera diagnostics to capture plasma evolution in the visible spectrum in real-time. These cameras have been instrumental in understanding plasma phenomena (Kirk2006), providing statistical insights into plasma turbulence (Walkden2022) and disruptions (Ham2022).
Building on our previous work (Gopakumar_2024), we apply the CP framework to an FNO trained to forecast plasma evolution from camera imagery. The model takes 10 consecutive camera frames as input and predicts the subsequent 10 frames. Using the AER nonconformity score, we demonstrate that CP provides statistically valid error bars for these predictions. Details on the camera data, FNO architecture, and training are available in (Gopakumar_2024).
3.8.1 Exchangeability and Coverage Validation
Although the FNO predicts plasma evolution over an entire shot duration, we structure the problem as an initial value problem where each forecast depends solely on its initial 10 frames. As illustrated in Figure 20, this allows us to treat each input-output pair as exchangeable: the model predicts the entire spatio-temporal output tensor simultaneously rather than sequentially, preserving the temporal structure.
The validity of CP for this application rests on three key properties: (1) only the calibration dataset and prediction $X_{n+1}$ must be exchangeable, (2) we predict error bars over the complete time interval simultaneously, not autoregressively, and (3) we assume minimal distributional shift between calibration and prediction shots. This final assumption requires that calibration and prediction shots exhibit similar plasma discharge profiles and device conditions.
To assess sensitivity to exchangeability violations, we compare coverage across shots with similar versus dissimilar plasma profiles relative to the calibration set. Figure 21 demonstrates that predictions on shots with substantially different characteristics suffer from insufficient coverage, highlighting the importance of the exchangeability assumption across the calibration and prediciton set. The calibration shot is indicative of a L-mode of confinement within the plasma, where as the inexchangeable prediciton shot shows H-mode of confinement characterising a vastly different physics regime (Howlett_2023). When this assumption holds, CP provides exact coverage as shown in Appendix I. Figure 22 visualises the calibrated error bars for plasma evolution forecasts across the tokamak central solenoid, demonstrating 50% coverage ($\alpha=0.5$).
@@ 3.8.1 Exchangeability and Coverage Validation
Figure 21: Coverage comparison for exchangeable (similar plasma profiles) versus non-exchangeable (dissimilar plasma profiles) prediction shots relative to the calibration set. Violations of exchangeability lead to insufficient coverage.
@@
Figure 22: Camera observations (top), FNO predictions (middle), and width of prediction intervals obtained using CP with $\alpha=0.5$ corresponding to 50% coverage (bottom).